OpenStudy (anonymous):

The diameter of the circle in the problem is 28 ft. Which is approximately equal to the area of the circle? Use 22 over 7 for pi. 88 sq ft. 308 sq ft. 616sq ft. 784 sq ft.

OpenStudy (anonymous):

The radius is 14 ft. Thus, the area is pi*r^2 = 22/7*14*14 = 616